# HF radar estimation of ocean wave parameters: second-order Doppler spectrum versus Bragg wave modulation approach

### v2

@@ 4.2 Global comparison with WWIII
Figure 12 shows the maps obtained for the statistical indices using the two methods, with the calibration factor $0.58$ applied to Barrick’s method estimate of $H_{s}$. The values considered to be irrelevant have been eliminated and shown in white. These are the radar cells for which the time coverage is lower than $30\%$ or the Pearson correlation coefficient is lower than $0.3$. The map confirms that Barrick’s method is more accurate than the Modulation method in estimating the $H_{s}$, with a RMSD of the order of $40$ cm in a large central zone (and probably better if the inaccuracy of WWIII itself is accounted for) out to about $30$ km range. The Pearson correlation coefficient remains larger than about $0.8$ in the same area. The Modulation method has in general a larger RMSD (of the order of $60$ cm at best) and smaller Pearson correlation coefficient (less than $0.7$). However, it provides coarse estimates with a high temporal coverage on a large distance. The same conclusions apply to the peak wave frequency (Figure 13) but with smaller values of the Pearson correlation coefficient for the two methods. Note that the modulation method is strongly affected around some ranges (number 15, 25-30 corresponding to distances of about 22 and 37-45 km) by the power grid artifact mentioned in Figure 6. However, relevant estimates are obtained in the far central zone and will now be discussed in more details.
@@
4.3 Estimation in the far range
As mentioned above, it is interesting to examine the robustness of the two methods with respect to the range. To do this, we calculated the statistical performance indices along two distinct radials. The first one is the azimuth $78$ of the CANDHIS buoy, which is marked with a thick black line in Figure 12; the second is azimuth $40$, which is directed towards the open sea and is marked with a thick magenta line. Figure 14 shows the evolution of the RMSD, Pearson correlation coefficient and time coverage along these two azimuths for the Barrick’s and Modulation method, where WWIII is again taken as the reference. As can be seen, the Modulation method remains relevant at long range in the open sea, with a Pearson correlation coefficient greater than $0.5$ even though the RMSD becomes large. This is visually confirmed by the time series of $H_{s}$ and $f_{p}$ shown in Figure 15 at the cell range $30$ ($45$ km). While Barrick’s method yields only sparse estimates, the Modulation method is able to follow the trends of the wave parameters, albeit with sometimes large errors. In addition, Barrick’s method fails in estimating the small or moderate values of SWH (less than about $1$ m) while the Modulation method remains sensitive to the variations of the SWH even for small values. In order to understand this difference, we compared the Doppler spectrum and the Modulation spectrum obtained over on hour at this range cell (Figure 16). While the second-order Doppler spectrum is no longer distinguishable, the amplitude modulation of the analytical signal is still visible and the corresponding power spectrum has a clear peak at a dominant frequency. This explains why this last method can provide relevant (if not accurate) estimates of the wave parameters out to a farther distance than Barrick’s method.
@@ 4.3 Estimation in the far range
Figure 14: Statistical performance indicators for Barrick’s method (blue lines) and the Modulation method (red lines). The black lines represent the RMS values of $H_{s}$ and $f_{p}$ according to the WWIII model at each range cell. The thick lines correspond to the azimuth $78$ (CANDHIS) while the thin line correspond to the azimuth $40$ (open sea).
@@
Figure 15: Time series of $H_{s}$ (top) and $f_{p}$ (bottom) in the long range ($45$ km) along the azimuth $40$ (magenta line of Figure 12) as estimated by the WWIII model (red lines), Barrick’s method (blue symbols) or the Modulation method (magenta symbols).
Figure 16: a) Example of a Doppler spectrum obtained over one hour in the long range ($45km$); b) Visualization of the amplitude modulation through the absolute analytical signal; c) Power spectrum of the absolute analytical signal.
@@ 4.4 Discussion
Barrick’s method, which is based on the second-order Doppler spectrum, has well-known limitations, which we summarize below. The roughness parameter $K_{0}H_{s}$, where $K_{0}$ is the electromagnetic wave number, should not be too large for otherwise the perturbation expansions in the scattering orders would break down. This is clearly satisfied in the present case since at $16.15$ MHz with the local wave conditions ($H_{s}<4$ m), we have $K_{0}H_{s}<1.2$ Another condition for the validity of the asymptotic formulae 5 and 7 (see Barrick, 1977a ) is that the peak wave frequency of the ocean wave spectrum be much lower than the Bragg frequency, that is $f_{p}<<f_{B}$, a condition that can be violated at low sea states. This explains why the quality of the estimation deteriorates as the mean frequency increases, as has already been mentioned in Section 4.1. In addition, the signal-to-noise ratio of the second-order Doppler spectrum increases with the sea state and may be insufficient in the case of a calm sea. For these reasons, Barrick’s method is not adapted for measuring small values of $H_{s}$ nor large values of $f_{p}$, a situation which in general corresponds to low wind speed. Beyond these conditions, Barrick’s approach is obviously limited in range because it requires the Doppler spectrum side lobes to be larger than the noise level. As a result, the estimation of wave parameters also has a shorter range than the estimation of current, which relies on the sole first-order Bragg peak. The limitations of the Modulation method are less obvious to establish. However, by construction this method also requires a clear separation of scales between the Bragg wave and the dominant ocean waves, which again excludes low sea states and peak wave frequencies near the Bragg frequency ($2.5$ seconds). The heuristic justification of the long-wave modulation is limited to small long wave steepness but does not a priori impose any condition on the roughness parameter $K_{0}H_{s}$. It is therefore expected to be robust at large wave heights. As it uses the full received voltage and not only its second-order part, the method is also expected to be able to cover a greater range than the Doppler-based method, as long as the slow amplitude modulation is visible. Since the latter is more pronounced when longer waves are more developed and have larger height and periods, the modulation method is expected to be favoured by high winds and to work better in the open sea than in coastal aeras. This is partly confirmed by our experimental study. However, some experimental artifacts (which are specific to our installation) such as the power grid instability, prevent us from using the full potential of the method and the full range of distances.
@@ 4.5 Comparison with the literature
A number of studies have been published in the literature on the estimation of peak wave heights and frequency from a single HFR station and a single radar frequency. Some of them reported systematic comparisons with in situ instruments over a significantly long period (at least several months) using either phased array (Wyatt et al.,, 2003; Ramos et al.,, 2009; Gomez et al.,, 2015; Lopez et al.,, 2016; Lopez and Conley,, 2019; Alattabi et al.,, 2019; Al-Attabi et al.,, 2021) or compact radar systems (Roarty et al.,, 2015; Zhou and Wen,, 2015; Long et al.,, 2011; Atan et al.,, 2016; Saviano et al.,, 2020; Basañez et al.,, 2020). All of these methods are based on the exploitation of the second-order Doppler spectrum using either Barrick’s original methods, the full wave spectrum inversion method, or empirical methods developed by the radar manufacturers. The performance obtained varies considerably depending on the geographical location, the radar frequency, the measurement range, the antenna array, the inversion method chosen, the data post-processing choices, etc. However, two general trends can be identified within this diversity of results. First, the estimation of $H_{s}$ with phased array systems is generally slightly better than that with compact systems, with high Pearson correlation coefficients ($>0.9$) and RMSD often in the order of 30-40 cm and less than about 60 cm in all studies. The observed RMSD with compact systems are generally larger, in the order of 40-80 cm, with a large scatter of values and in some cases very large RMSD ($>1$ m); the associated Pearson correlation coefficients are generally smaller, mostly in the range of 0.5-0.9, except for the study Atan et al., (2016), which reports exceptionally good performances of 18-29 cm RMSD and 0.78-0.86 Pearson correlation coefficient. However, the decision to eliminate the $>25$% outliers (when compared with buoy measurement) before calculating the statistics is questionable. The second point is that the estimation of the peak frequency is less systematic and generally of lower quality than that of $H_{s}$, with Pearson correlation coefficients of the order of 0.5-0.7.
@@ 5 Conclusion
In this work, we have tested the accuracy of the classical Barrick’s method for estimating the main sea state parameters based on weighted integrals of the second-order backscattered Doppler spectrum. For this purpose, we analyzed a one-year HFR data set from a monostatic station in the Toulon region. The results presented in this study have been compared with the data collected by the CANDHIS buoy located near the island of Porquerolles as well as hourly data simulated from the high-resolution WWIII model. For the classical Barrick’s method, the results obtained for the SWH have a good accuracy compared to the time series of the CANDHIS buoy, with a very high correlation and a low RMSD value between both of them (close to $90\%$ and around $0.3$ m, respectively, at $30$ km from the radar). A scaling factor of $0.58$ was found to be necessary in order to adjust the statistics of the two data sets and to further reduce the RMSD. With regard to the estimation of the peak wave frequency, this method has some shortcomings in capturing high values, due to insufficient scale separation between the long wave and the Bragg wave.
In addition, a new method, called the “Bragg Wave Modulation Method” (the “Modulation Method” for short), was also introduced and evaluated together with Barrick’s method. This technique uses the amplitude modulation of the received voltage time series and does not assume a decomposition of the signal in perturbation orders. It was found to be much less accurate than Barrick’s method at this stage but has been found to be promising for estimating wave characteristics at greater distances from the radars. A further limitation of this new method, at least for the $H_{s}$, is that it requires a preliminary calibration procedure with some ground truth measurements. This is not the case with Barrick’s method, which is based on the ratio of the first and second-order Doppler spectra and is therefore insensitive to absolute calibration (although a rather universal scaling factor improves its accuracy). The “Modulation Method” will be pursued and hopefully improved in future works.

### v3

@@ 4.2 Global comparison with WWIII
Figure 12 shows the maps obtained for the statistical indices using the two methods, with the calibration factor $0.58$ applied to Barrick’s method estimate of $H_{s}$. The values considered to be irrelevant have been eliminated and shown in white. These are the radar cells for which the time coverage is lower than $30\%$ or the Pearson correlation coefficient is lower than $0.3$. The map confirms that Barrick’s method is more accurate than the Modulation method in estimating the $H_{s}$, with a RMSD of the order of $40$ cm in a large central zone (and probably better if the inaccuracy of WWIII itself is accounted for) out to about $30$ km range. The Pearson correlation coefficient remains larger than about $0.8$ in the same area. The Modulation method has in general a larger RMSD (of the order of $60$ cm at best) and smaller Pearson correlation coefficient (less than $0.7$). Because no constraint was imposed at this stage in the Modulation method (such as e.g. threshold conditions regarding certain quantities), it has by construction a higher spatial and temporal coverage than the Doppler-based method. The same conclusions apply to the peak wave frequency (figure 13) but with smaller values of the Pearson correlation coefficient for the two methods. Note that the modulation method is strongly affected around some ranges (number 15, 25-30 corresponding to distances of about 22 and 37-45 km) by the power grid artifact mentioned in figure 6. However, relevant estimates are obtained in the far central zone and will now be discussed in more details.
@@
4.3 Statistical performances in the long range
We now examine the robustness of the two methods with respect to the range. To do this, we calculated the statistical performance indices along two distinct radials. The first one is the azimuth $78$ of the CANDHIS buoy, which is marked with a thick black line in figure 12; the second is azimuth $40$, which is directed towards the open sea and is marked with a thick magenta line. Figure 14 shows the evolution of the RMSD, Pearson correlation coefficient and time coverage along these two azimuths for the Barrick’s and Modulation method, where WWIII is again taken as the reference. As expected, the RMS error for the $H_{s}$ is increasing with range while the Pearson correlation coefficient is decreasing. However, the RMS errors obtained with the Modulation method are much larger than those obtained with Barrick’s method and the Pearson correlation coefficient is much smaller. Taking as a maximum acceptable error the mean annual value of $H_{s}$ calculated at each range with the WWIII model, we see that the $H_{s}$ estimates obtained with the modulation method are relevant up to a distance of about 22 km along the azimuth 78 (CANDHIS direction) and 46 km along the azimuth $40$ (open sea). In spite of this increasing RMS error, the Pearson correlation coefficient obtained with the Modulation method remains meaningful (say, greater than $0.5$) over large distances, especially in the open sea. Regarding the $f_{p}$ estimate, the method performances are less contrasted, especially in the open sea. Their RMS errors show little dependence upon range and remain much smaller than the mean parameter value. The Pearson correlation remains larger than about $0.6$ up to a distance of $35$ km. To better evaluate the quality of estimation in the long range, we plotted the time series of the estimated $H_{s}$ and $f_{p}$ at the cell range $32$ ($48$ km) (figure 15). The $H_{s}$ estimates with the Modulation method follow the overall trend of variation of the WWIII model but suffer from important errors; the $f_{p}$ estimates are very weakly correlated to the WWIII data and unable to follow their full variability. On the contrary Barrick’s method yields sparse but relevant estimates, especially for the $H_{s}$. This is confirmed by the scatter plots in figure 16, which shows large scatter of estimates with the Modulation method while the rare estimates obtained with Barrick’s method are well aligned with the WWIII values. It must be noted, however, that the selection tests that have been employed to identify the transition between the first- and second-order Doppler spectrum have made this last method more reliable. Again, it must be mentioned that we did not impose any test criterion to the Modulation method while evaluating the wave parameters. This explains at the same time its much higher time coverage and its much higher level of error.
@@ 4.3 Statistical performances in the long range
Figure 14: Statistical performance indicators for Barrick’s method (blue lines) and the Modulation method (red lines). The black lines represent the mean values of $H_{s}$ and $f_{p}$ according to the WWIII model at each range cell. The thick lines correspond to the azimuth $78$ (CANDHIS) while the thin line correspond to the azimuth $40$ (open sea).
@@
Figure 15: Time series of $H_{s}$ (top) and $f_{p}$ (bottom) in the long range ($48$ km) along the azimuth $40$ (magenta line of figure 12) as estimated by the WWIII model (red lines), Barrick’s method (blue symbols) or the Modulation method (magenta symbols).
Figure 16: Scatter plots of $H_{s}$ and $f_{p}$ in the long range ($48$ km) along the azimuth $40$ (magenta line of figure 12) as estimated with Barrick’s method (top) and the Modulation method (bottom) and compared with the WWIII model.
@@ 4.4 Discussion
Barrick’s method, which is based on the second-order Doppler spectrum, has well-known limitations, which we summarize below. The roughness parameter $K_{0}H_{s}$, where $K_{0}$ is the electromagnetic wave number, should not be too large for otherwise the perturbation expansions in the scattering orders would break down. This is clearly satisfied in the present case since at $16.15$ MHz with the local wave conditions ($H_{s}<4$ m), we have $K_{0}H_{s}<1.2$ Another condition for the validity of the asymptotic formulae 5 and 7 (see Barrick, 1977a ) is that the peak wave frequency of the ocean wave spectrum be much lower than the Bragg frequency, that is $f_{p}<<f_{B}$, a condition that can be violated at low sea states. This explains why the quality of the estimation deteriorates as the mean frequency increases, as has already been mentioned in Section 4.1. In addition, the signal-to-noise ratio of the second-order Doppler spectrum increases with the sea state and may be insufficient in the case of a calm sea. For these reasons, Barrick’s method is not adapted for measuring small values of $H_{s}$ nor large values of $f_{p}$, a situation which in general corresponds to low wind speed. Beyond these conditions, Barrick’s approach is obviously limited in range because it requires the Doppler spectrum side lobes to be larger than the noise level. As a result, the estimation of wave parameters also has a shorter range than the estimation of current, which relies on the sole first-order Bragg peak. The limitations of the Modulation method are less obvious to establish. However, by construction this method also requires a clear separation of scales between the Bragg wave and the dominant ocean waves, which again excludes low sea states and peak wave frequencies near the Bragg frequency ($2.5$ seconds). The heuristic justification of the long-wave modulation is limited to small long wave steepness but does not a priori impose any condition on the roughness parameter $K_{0}H_{s}$. It is therefore expected to be robust at large wave heights. As it uses the full received voltage and not only its second-order part, the method is also expected to be able to cover a greater range than the Doppler-based method, as long as the slow amplitude modulation is visible. Since the latter is more pronounced when longer waves are more developed and have larger height and periods, the modulation method is expected to be favoured by high winds and to work better in the open sea than in coastal aeras. This is partly confirmed by our experimental study. However, at this early stage, the modulation method is not able to challenge Barrick’s method in any respect. Although it produces meaninful estimates at near and middle range and show some correlation with the WWIII model at long range, it is far too inaccurate at the moment to be a useful complement to Barrick’s method. In addition, it suffers from some experimental artifacts (which are specific to our installation) such as the power grid instability, which prevent us from using the method over the full range of distances.
@@ 4.5 Comparison with the literature
A number of studies have been published in the literature on the estimation of peak wave heights and frequency from single or multiple HFR stations and a single radar frequency. Some of them reported systematic comparisons with in situ instruments over a significantly long period (at least several months) using either phased array (Wyatt et al.,, 2003; Ramos et al.,, 2009; Gomez et al.,, 2015; Lopez et al.,, 2016; Lopez and Conley,, 2019; Alattabi et al.,, 2019; Al-Attabi et al.,, 2021) or compact radar systems (Roarty et al.,, 2015; Zhou and Wen,, 2015; Long et al.,, 2011; Atan et al.,, 2016; Saviano et al.,, 2020; Basañez et al.,, 2020). All of these methods are based on the exploitation of the second-order Doppler spectrum using either Barrick’s original methods, the full wave spectrum inversion method, or empirical methods developed by the radar manufacturers. The performance obtained varies considerably depending on the geographical location, the radar frequency, the number of stations, the measurement range, the antenna array, the inversion method chosen, the data post-processing choices, etc. However, two general trends can be identified within this diversity of results. First, the estimation of $H_{s}$ with phased array systems is generally slightly better than that with compact systems, with high Pearson correlation coefficients ($>0.9$) and RMSD often in the order of 30-40 cm and less than about 60 cm in all studies. The observed RMSD with compact systems are generally larger, in the order of 40-80 cm, with a large scatter of values and in some cases very large RMSD ($>1$ m); the associated Pearson correlation coefficients are generally smaller, mostly in the range of 0.5-0.9, except for the study Atan et al., (2016), which reports exceptionally good performances of 18-29 cm RMSD and 0.78-0.86 Pearson correlation coefficient. However, the decision to eliminate the $>25$% outliers (when compared with buoy measurement) before calculating the statistics is questionable. The second point is that the estimation of the peak frequency is less systematic and generally of lower quality than that of $H_{s}$, with Pearson correlation coefficients of the order of 0.5-0.7.
@@ 5 Conclusion
In this work, we have tested the accuracy of the classical Barrick’s method for estimating the main sea state parameters based on weighted integrals of the second-order backscattered Doppler spectrum. For this purpose, we analyzed a one-year HFR data set from a monostatic station in the Toulon region. The results presented in this study have been compared with the data collected by the CANDHIS buoy located near the island of Porquerolles as well as hourly data simulated from the high-resolution WWIII model. For the classical Barrick’s method, the results obtained for the SWH have a good accuracy compared to the time series of the CANDHIS buoy, with a very high correlation and a low RMSD value between both of them (close to $90\%$ and around $0.3$ m, respectively, at $32$ km from the radar). A scaling factor of $0.58$ was found to be necessary in order to adjust the statistics of the two data sets and to further reduce the RMSD. With regard to the estimation of the peak wave frequency, this method has some shortcomings in capturing high values, due to insufficient scale separation between the long wave and the Bragg wave.
In addition, a new method, called the “Bragg Wave Modulation Method” (the “Modulation Method” for short), was also introduced and evaluated together with Barrick’s method. This technique uses the amplitude modulation of the received voltage time series and does not assume a decomposition of the signal in perturbation orders. It was found to be much less accurate than Barrick’s method at this stage and needs to be complemented with further quality tests. A further limitation of this new method, at least for the $H_{s}$, is that it requires a preliminary calibration procedure with some ground truth measurements. This is not the case with Barrick’s method, which is based on the ratio of the first and second-order Doppler spectra and is therefore insensitive to absolute calibration (although a rather universal scaling factor improves its accuracy). The “Modulation Method” will be pursued and hopefully improved in future works.